# SIV-Bench: A Video Benchmark for Social Interaction Understanding and Reasoning

### v1

@@ 2.3 Comparison with Previous Datasets & Benchmarks
Currently, most general video understanding benchmarks (like VideoVista [25] and MVBench [24]) lack a focus on social interaction; others (like MLVU [61] and Video-Bench [32]), while touching upon it, do not fully cover social relations. Even Social-IQ 2.0 [51], which concentrates on this area, has limitations in task diversity and dynamic reasoning. In contrast, SIV-Bench is built on original data, combines manual and automatic annotations, and is one of the few benchmarks to provide subtitle and audio information, thereby supporting richer multimodal social reasoning.
Table 1: Comparison of various benchmarks. It includes several aspects: total number of items (#Items, representing the number of videos, dialogues, images, etc.), number of QA pairs (#QAs), annotation method (Anno., M/A means manually/automatic manner), and the types of tasks included (Task Types: SU for Scene Understanding, SR for State Reasoning, DP for Dynamics Prediction). Note that Task Types here refer to general task categories, which include both social and physical scenarios. The table also shows whether each benchmark includes Multi-Person Interaction, covers Various Relations, is based on newly collected data (Original Collection), and provides subtitle/audio information (S.A.)
Benchmark #Items #QAs Anno. Task Types Multi-Person Various Original S.A. SU SR DP Interaction Relations Collection Social Relation Inference DialogRE [58] 1,788 - M - PIPA [45] 37,107 - M - ViSR [26] 8,000 - M General Video Understanding and Reasoning VideoMME [13] 900 2,700 M MLVU [61] 1,730 3,102 M VideoVista [25] 894 24,906 A Social-IQ 2.0 [51] 1,000 6,000 M Perception Test [34] 11,620 38,000 M MVBench [24] 3,641 4,000 A Video-Bench [32] 5,917 17,036 A&M EgoSchema [28] 5,063 5,063 A&M SIV-Bench 2,792 8,728 A&M
@@ 3.1 Settings
We evaluate a diverse set of closed- and open-source MLLMs on SIV-Bench, including Gemini-2.0/2.5-Flash [15, 14], Gemini-2.5-Pro [9], GPT-4o [20], o4-mini [33], Qwen2.5-VL-7B/72B-Instruct [3], mPLUG-Owl3 [57], InternVL3-8B/78B [63], LLaVA-OneVision [23], and LLaVA-Video [59]. Evaluations are conducted using VLMEvalKit [10], with models generating responses under their default settings. Models capable of direct video processing (e.g., Gemini) are fed raw videos, while image-only models (e.g., o4-mini) receive 16 uniformly sampled frames as input.
We use a standardized prompt (Figure 18) across all models, instructing them to return both the answer letter (e.g., ‘A.’) and the full answer text. However, we observe variations in model compliance: for instance, some models (e.g., InternVL-8B and mPLUG-Owl3) occasionally tend to output only the letter, whereas others (e.g., LLaVA-Video) may provide only the text. To robustly handle these diverse output formats and ensure accurate answer parsing, we implement a two-stage matching procedure. First, we check the model’s output for a correctly formatted option letter. If absent, we use a similarity-based comparison between the model’s raw output and the full text of all answer options to identify its selection. Model performance is then measured by the accuracy of these inferred responses, as shown in Table 2. The rest of this section provides a detailed analysis.
3.2 Main Results
Table 2: Evaluation results of MLLMs on SIV-Bench. Scores are reported for SSU, SSR, SDP, and Overall performance, detailing the impact of subtitle conditions: ‘origin,’ ‘+sub’ (added transcribed dialogue), and ‘-sub’ (subtitles removed). All scores are presented as accuracy percentages (%).
Models Params Social Scene Understanding Social State Reasoning Social Dynamics Prediction Overall origin + sub - sub origin + sub - sub origin + sub - sub origin + sub - sub Open-source MLLMs mPLUG-Owl3 7B 66.32 66.21 66.46 62.36 62.19 61.33 65.19 66.30 65.22 63.79 64.01 63.22 LLaVA-OneVision 7B 62.00 62.13 62.55 63.72 64.78 61.66 64.87 65.32 62.14 63.73 64.40 62.14 LLaVA-Video 7B 68.89 69.13 69.10 62.08 61.37 60.09 63.50 63.84 62.46 63.18 63.75 61.66 Qwen2.5-VL-7B-Instruct 7B 69.51 69.30 68.89 62.65 61.84 61.04 64.18 64.89 63.90 65.01 65.13 63.53 InternVL3-8B 8B 73.02 72.58 72.81 62.72 63.06 61.20 65.33 65.95 65.25 66.14 66.28 65.35 Qwen2.5-VL-72B-Instruct 72B 84.83 85.15 83.46 70.23 70.54 69.58 74.88 74.49 74.11 74.25 74.77 73.54 InternVL3-78B 78B 82.16 83.54 82.35 67.45 67.91 66.87 73.14 73.46 72.21 72.16 72.70 71.56 Closed-source MLLMs o4-mini - 86.77 86.90 86.33 66.87 67.25 66.10 72.86 72.45 72.11 72.30 72.43 71.59 GPT-4o - 86.94 87.34 86.29 68.28 68.58 67.73 74.82 75.80 74.05 73.76 74.29 73.12 Gemini-2.0-Flash - 86.54 86.35 86.46 67.73 68.07 66.49 73.14 73.79 72.42 72.75 73.27 71.65 Gemini-2.5-Flash - 88.56 88.84 87.32 68.12 69.09 67.25 74.67 74.97 73.05 73.67 73.82 72.53 Gemini-2.5-Pro - 90.67 90.88 90.59 71.44 71.78 70.20 75.28 75.96 74.27 76.03 76.50 75.14
Overall Performance. The evaluation of various models on SIV-Bench reveals distinct performance tiers, with closed-source models generally setting the current state-of-the-art. Notably, Gemini-2.5-Pro emerges as the top-performing model, achieving the highest overall accuracy of 76.50% when provided with subtitles. Other prominent closed-source models, such as GPT-4o and Gemini-2.5-Flash, also demonstrate strong overall performance, consistently scoring above 73%. These results indicate the advanced capabilities of leading proprietary systems.
Table 3: Average relative performance change (%) of all models when adding (‘+sub’) or removing (‘-sub’) subtitles, compared to the ‘origin’.
Condition SSU (%) SSR (%) SDP (%) Overall (%) + sub 0.178 0.234 0.482 0.447 - sub -0.300 -1.176 -0.878 -0.878
Following these, open-source Multi-modal Large Language Models (MLLMs) show considerable capabilities, although a performance gap still exists. Among them, models with a larger number of parameters exhibit significantly better results; for example, Qwen2.5-VL-72B-Instruct achieves an overall score of 74.77% with subtitles, markedly outperforming its 7B counterpart (65.13% with subtitles), even superior to the closed-source models except for Gemini-2.5-Pro. Similarly, InternVL-78B performs substantially better than InternVL-8B. While these larger open-source models demonstrate strong potential and achieve commendable overall scores, there remains room for improvement to match the leading closed-source alternatives.
Subtitle Influence. Our analysis of subtitle influence (average relative changes shown in Table 3) shows that while explicit subtitles benefit MLLMs, their impact varies across tasks. SSU, which relies on direct visual cues like actions and environments, is least affected by subtitle changes, with minimal performance shifts when subtitles are added (+sub, +0.207%) or removed (-sub, -0.333%), suggesting that visual perception alone provides sufficient grounding for this task. In contrast, the more inferential tasks—SSR and SDP—show greater sensitivity to linguistic input. Transcribed dialogue (+sub) improves SDP (+0.654%) by clarifying narrative flow and communicative intent, while removing all subtitles (-sub) most impacts SSR (-1.806%), where understanding social states becomes harder without dialogue. These findings indicate that the linguistic cues from dialogue play an important role in understanding and reasoning about multi-modal social interactions.
Task-Specific Performance. SSU, focused on recognizing visible elements, is typically the easiest, with models achieving their highest scores here. However, this isn’t always the case—for instance, LLaVA-OneVision sometimes performs better on SSR or SDP. In SSU, stronger models—especially large-scale ones, whether closed-source or high-parameter open-source—consistently outperform smaller models. This performance gap, particularly pronounced in this foundational visual task, likely reflects the superior perceptual capabilities of larger models due to greater capacity and more extensive training [1]. Following SSU, SDP tends to be the next most tractable, as it involves reasoning about interactions and predicting outcomes. SSR, which requires inferring unobservable social states like emotions and intentions, remains the most difficult. Even the top models score lowest here, underscoring the challenge of capturing implicit psychological dimensions.
Figure 6: Radar chart illustrating the comparative performance of evaluated MLLMs across the 10 fine-grained sub-tasks within SIV-Bench.
@@ 3.2 Main Results
To further break down model capabilities, a radar chart (Figure 6) visualizes performance across 10 fine-grained tasks in original videos (with ‘+sub’ and ‘-sub’ detailed in Appendix C.2). Stronger models, particularly closed-source ones like the Gemini series and GPT-4o, form the outer ring, clearly outperforming smaller open-source models. In SSU, the performance gap is especially evident in Action Recognition (AR) and Facial Expression Recognition (FER), reflecting the stronger models’ ability to capture subtle visual cues. SSR presents greater challenges: while models perform moderately on Intent Inference (II) and Emotion Inference (EI), they struggle with more abstract tasks like Relation Inference (RI), which often mark the lowest points on the radar. This illustrates the difficulty of moving from surface-level cues to deeper social reasoning. In SDP, performance improves slightly, likely because tasks such as Factual Prediction (FP) and Counterfactual Prediction (CP) tap into social commonsense learned from language data. Most models—except a few like LLaVA-OneVision—perform better on CP than FP, possibly due to the more explicit reasoning cues provided by hypothetical framing. Some failure cases are illustrated in Figure 19, 20, 21.
3.3 Failure Pattern Analysis of Relation Inference
Accurate Relation Inference (RI) is crucial for understanding social interactions. The difficulty lies in synthesizing diverse cues—scene context, human appearance, verbal and non-verbal expressions, physical proximity, and interaction dynamics—all of which contribute to subtle and implicit relational patterns. Our fine-grained analysis (Figure 6) highlights RI as a consistent performance bottleneck. Given its significance and the persistent challenges across models, we conduct a deeper qualitative analysis to examine common failure patterns:
Lack of Primary-Secondary Relation Differentiation. In multi-relational scenarios, MLLMs sometimes identify a valid but secondary relationship instead of the most salient one. For instance, in a video where employees celebrate their boss’s birthday, the model misclassifies the situation as <colleagues>, overlooking the underlying hierarchical relationship that defines the interaction. Similarly, when a coach angrily reprimands players for their poor performance, the model incorrectly assigns the relationship as <teammates>, failing to capture the coach-athlete authority structure that governs the scene. These errors suggest a deficiency in relational prioritization and in the ability to infer which social bond is most significant within a complex interpersonal context.
Scene and Human-Induced Misleading Cues. MLLMs are easily misled by both scene settings and human behaviors. They may over-rely on environmental cues, assigning stereotypical roles based on location without validating actual interactions. For instance, in a classroom setting where two travel agency employees co-present, it incorrectly classifies them as <teacher-student> instead of <colleagues>. Similarly, human language can mislead the model: a couple jokingly referring to each other as "best brother" is misclassified as <siblings>, despite strong visual evidence—such as different ethnicities and intimate gestures—indicating a romantic relationship.
Deficiency in Commonsense Reasoning. In several instances, the model’s decisions are inconsistent with basic commonsense knowledge. Two <friends> wearing different sports team uniforms but showing friendly behavior are misclassified as <teammates>, disregarding the fact that such uniforms typically signify competition rather than collaboration. Similarly, a store owner photographing a bustling crowd is mistaken for a colleague of the customers, whereas commonsense suggests a <transactional> relationship. These mistakes highlight the model’s limited ability to integrate external knowledge about social conventions and situational functions, emphasizing a shortfall in commonsense reasoning and real-world social understanding.
Ignoring Key Perceptual Details. Although MLLMs can capture many surface features, they will overlook subtle but critical details necessary for correct role recognition. Age differences and facial features crucial to distinguishing between <siblings> and <parent-child> relationships are sometimes ignored. Additionally, important textual clues embedded in subtitles are sometimes overlooked, such as the line "POV: you work with your coworker," which results in <colleague> not being correctly assigned. These errors suggest insufficient fine-grained perception, selective attention to critical multimodal information, and a failure to integrate minor yet decisive cues into final role judgments.
@@ 4 Related Works
Video Benchmarks for MLLMs. The evaluation of rapidly advancing Multimodal Large Language Models (MLLMs) in the video domain relies on diverse benchmarks that collectively incorporate a rich set of features for comprehensive model assessment. Many, such as MLVU [61], video-MME [13], MVbench [24], MMBench-Video [11], and Video-Bench [32], offer multiple tasks—ranging from question answering and summarization to complex reasoning—across diverse video fields including movies, egocentric recordings, and general web content. This variety aims to assess a broad spectrum of capabilities, from foundational visual understanding and detail analysis (as targeted by Perception Test [34]) (explored by VideoVista [25]) to higher-level comprehension and decision-making. A significant trend within these evaluations is the robust focus on long-form video understanding, with benchmarks like CINEPILE [40], Egoschema [28], and LVbench [50] specifically challenging models on temporal reasoning and extended context retention, while others like TempCompass [27] provide specialized assessments of critical nuanced aspects such as detailed temporal dynamics. While these benchmarks thoroughly assess diverse video comprehension capabilities, SIV-Bench offers a more specialized focus, concentrating on the intricate understanding and reasoning required to interpret dynamic social interactions among multiple people.
Evaluating Social Intelligence in AI Systems. The pursuit of artificial social intelligence is inherently multifaceted, with research progressing from early dataset construction to contemporary evaluations of large language models (LLMs). Initial efforts typically offered valuable yet modality-constrained or component-specific resources, targeting areas such as action recognition (e.g., Kinetics [22], UCF101 [43]), facial expression and contextual emotion inference (e.g., AffectNet [31], MELD [35]), and social relation recognition from static images or text (e.g., PIPA [45], DialogRE [58]). As the field has evolved, benchmarks like Social-IQ 2.0 [51] have expanded the focus toward interpersonal reasoning and social commonsense. With the rise of LLMs, more specialized evaluations have emerged. ToMATO [41] probes Theory of Mind through role-play-based interactions. SOTOPIA [62] and its extension SOTOPIA-$\pi$ [49] introduce interactive environments for developing agentive social skills. EgoSocialArena [18] offers a first-person perspective benchmark targeting both cognitive and behavioral intelligence. Social Genome [29] provides multimodal, grounded traces of social reasoning from video data, while MM-SOC [21] evaluates MLLMs on multimodal social media content. Despite these advances, most existing benchmarks focus on isolated competencies. A comprehensive evaluation framework that holistically assesses MLLMs’ understanding of complex social interactions—spanning perception, social state reasoning, and dynamics prediction in real-world video contexts—remains lacking. SIV-Bench is introduced to fill this gap.

### v2

@@ 2.3 Comparison with Existing Benchmarks
Table 1 highlights the key differences between SIV-Bench and existing works. Prior datasets on social interaction, such as DialogRE (Yu et al., 2020) (text-only), PIPA (Sun et al., 2017) (image-only), and ViSR (Liu et al., 2019), are limited to relation recognition, single modalities, and lack task diversity and scalability due to manual annotation. Currently, most general video understanding benchmarks (like VideoVista (Li et al., 2024c) and MVBench (Li et al., 2024b)) lack a focus on social interaction; others (like MLVU (Zhou et al., 2024) and Video-Bench (Ning et al., 2023)), while touching upon it, do not fully cover social relations. Even Social-IQ 2.0 (Wilf et al., 2023), which concentrates on this area, has limitations in task diversity and dynamic reasoning. SIV-Bench is built on original data, combines manual and automatic annotations, and is one of the few benchmarks to provide subtitle and audio information, supporting richer multimodal social reasoning.
Table 2: Evaluation results of MLLMs on SIV-Bench. Accuracy (%) on SSU, SSR, SDP, and Overall under different subtitle settings (‘origin’, ‘+sub’, ‘-sub’). Statistical significance tests are reported in Appendix C.5.
Models Params Social Scene Understanding Social State Reasoning Social Dynamics Prediction Overall origin + sub - sub origin + sub - sub origin + sub - sub origin + sub - sub Open-source MLLMs mPLUG-Owl3 7B 46.11 45.94 46.34 39.78 39.50 38.13 44.30 46.08 44.35 42.06 42.42 41.15 LLaVA-OneVision 7B 39.20 39.41 40.08 41.95 43.65 38.66 43.79 44.51 39.42 41.97 43.04 39.42 LLaVA-Video 7B 50.22 50.61 50.56 39.33 38.19 36.14 41.60 42.14 39.94 41.09 42.00 38.66 Qwen2.5-VL-7B-Instruct 7B 51.22 50.88 50.22 40.24 38.94 37.66 42.69 43.82 42.24 44.02 44.21 41.65 InternVL3-8B 8B 56.83 56.13 56.50 40.35 40.90 37.92 44.53 45.52 44.40 45.82 46.05 44.56 Qwen2.5-VL-72B-Instruct 72B 75.73 76.24 73.54 52.25 52.75 51.21 59.02 58.40 57.78 58.80 59.63 57.66 InternVL3-78B 78B 71.46 73.66 71.76 51.65 52.39 50.14 55.77 56.28 54.25 55.46 56.32 54.50 Closed-source MLLMs o4-mini - 78.83 79.04 78.13 50.47 51.30 48.99 56.89 56.00 55.26 55.68 55.89 54.54 GPT-4o - 79.10 79.74 78.06 52.73 53.20 51.79 59.02 60.59 58.60 58.02 58.86 56.99 Gemini-2.0-Flash - 78.46 78.16 78.34 51.89 52.43 49.78 57.59 58.63 55.70 56.40 57.23 54.64 Gemini-2.5-Flash - 81.70 82.14 79.71 48.99 50.54 47.60 59.47 59.95 56.88 57.87 58.11 56.05 Gemini-2.5-Pro - 85.07 85.41 84.94 54.30 54.85 52.32 60.45 61.54 58.83 61.65 62.40 60.22
@@ 3.1 Experimental Setup
We evaluate a diverse set of closed- and open-source MLLMs on SIV-Bench, including Gemini-2.0/2.5-Flash (Google, 2025a, b), Gemini-2.5-Pro (Doshi, 2025), GPT-4o (Hurst et al., 2024), o4-mini (OpenAI, 2025), Qwen2.5-VL-7B/72B-Instruct (Bai et al., 2025), mPLUG-Owl3 (Ye et al., 2024), InternVL3-8B/78B (Zhu et al., 2025a), LLaVA-OneVision (Li et al., 2024a), and LLaVA-Video (Zhang et al., 2024). Evaluations are conducted using VLMEvalKit (Duan et al., 2024), which provides a unified evaluation interface across MLLMs. All models generate responses under their default inference settings to ensure a fair comparison.
We use a standardized prompt (Figure 18) across all models, asking for both an option letter (e.g., ‘A.’) and the full answer text. To parse outputs robustly, we first extract a valid option letter; if missing, we match the raw output to all options using text similarity. To enable scalable and model-agnostic evaluation across heterogeneous MLLMs, we cast all tasks into a unified multiple-choice interface and report accuracy as the primary metric.
3.2 Overall Performance
The results shown in Table 2 highlights a clear performance stratification primarily driven by model scale. While the proprietary Gemini-2.5-Pro achieves the state-of-the-art (62.40% w/ subtitles), large-scale open-source models demonstrate exceptional competitiveness; notably, Qwen2.5-VL-72B (59.63%) outperforms other leading proprietary systems like GPT-4o and Gemini-2.5-Flash (>58%). A significant capacity gap is evident, as larger models consistently dominate their smaller counterparts (e.g., Qwen-72B’s 59.63% vs. 7B’s 44.21%), suggesting that sufficient parameter count is a prerequisite for robust social reasoning. We verify that the observed performance gaps are statistically significant using paired bootstrap tests; full details are reported in Appendix C.5.
3.3 Decomposed Performance
Table 3: Relative performance change (%) against the baseline. The audio ablation (w/o Audio) is evaluated on Gemini-2.5-Flash, while subtitle variations (+Sub/-Sub) reflect the average impact across all models.
Condition SSU SSR SDP Overall
w/o Audio -0.35 -2.44 -2.33 -1.40
+ Subtitle +0.05 +0.28 +0.65 +0.15
- Subtitle -0.97 -2.07 -1.68 -0.90
Subtitle and Audio Influence. To assess the impact of different modalities, we analyze the relative performance changes compared to the original video baseline, as shown in Table 3. The audio ablation is conducted on Gemini-2.5-Flash, while subtitle variations represent the average results across all evaluated models. The results demonstrate a clear hierarchy of modality importance. Removing audio leads to the most substantial performance drop (-1.40%), particularly in reasoning-intensive tasks (SSR: -2.44%), highlighting the necessity of auditory cues for social interaction understanding and reasoning. For linguistic cues, the impact is notably asymmetric: while adding subtitles provides only marginal gains (+0.15%), their removal significantly hinders performance (-0.90%), especially in SSR and SDP. This suggests that while models may not always benefit from redundant text, the absence of critical textual context creates a significant bottleneck for complex social reasoning.
@@ 3.3 Decomposed Performance
Task Influence. SSU focuses on recognizing visible elements and is generally the easiest dimension, where models achieve their highest scores. This is not always the case, since LLaVA-OneVision sometimes performs better on SSR or SDP. In SSU, larger models, including closed-source models and high-parameter open-source models, consistently outperform smaller ones, which likely reflects stronger perceptual ability from greater capacity and broader training (Alabdulmohsin et al., 2022). After SSU, SDP is usually the next most tractable, while SSR remains the most difficult because it requires inferring latent social states such as emotions and intentions. Figure 6 further breaks down performance across 10 fine-grained tasks in original videos. Stronger closed-source models such as the Gemini and GPT-4o form the outer ring and outperform smaller open-source models. In SSU, the performance gap is most pronounced in Action Recognition (AR) and Facial Expression Recognition (FER), suggesting advantages in capturing subtle visual cues. SSR is more challenging. Models perform moderately on Intent Inference (II) and Emotion Inference (EI), but they struggle most with Relation Inference (RI), which is often the lowest point on the radar. In SDP, performance improves on Factual Prediction (FP) and Counterfactual Prediction (CP), which may reflect social commonsense acquired from language data. Most models perform better on CP than FP, possibly because hypothetical framing provides clearer reasoning cues. Representative failures are shown in Figure 19, 20, and 21.
Figure 6: Radar chart of MLLM performance across the 10 fine-grained SIV-Bench sub-tasks.
3.4 In-Depth Analysis of Relation Inference
Since RI emerges as a key bottleneck and provides the relational grounding needed to interpret many higher-level social states, we take a closer look at RI and characterize its error structure quantitatively. Figure 7 aggregates RI predictions across the evaluated models into a confusion matrix, which reveals structured error clusters rather than random noise: a prominent pattern is Authority–Equality confusion, where hierarchical relations such as Boss–Employee and Coach–Player are frequently predicted as their egalitarian counterparts (Colleagues and Teammates). Errors also concentrate within relation families. The red dashed boundaries delineate the four foundational relational models, and off-diagonal clusters around these boundaries highlight systematic confusions between conceptually adjacent relation groups.
Figure 7: Aggregated Confusion Matrix for the Relation Inference (RI) task across all models. Red dashed lines delineate the four foundational relational models.
Qualitatively, we observe four common failure modes. (1) Models fail to differentiate primary vs. secondary relations in multi-relational scenarios, predicting a plausible but less salient relation. (2) Models are misled by scene- and human-induced cues, over-relying on stereotypical settings or surface language without validating the actual interaction (e.g., classroom $\rightarrow$ teacher–student). (3) Some errors reflect deficient commonsense social reasoning, where predictions contradict basic social conventions (e.g., two <friends> wearing different sports team uniforms but showing friendly behavior are misclassified as <teammates>, disregarding the fact that such uniforms typically signify competition). (4) Models ignore key perceptual details (e.g., age/role cues) that are decisive for disambiguating closely related relations. More failure cases are provided in Appendix C.2.
3.5 In-depth Analysis of Reasoning Process
To diagnose the cognitive boundaries of current MLLMs, we establish SIV-Bench-Hard, a subset of 200 QAs where models collectively demonstrate the highest error rates. This focused selection enables concentrated failure-mode analysis and facilitates a reliable human baseline, which is prohibitively expensive at full scale. We recruit three independent crowdworkers, distinct from the original annotators, to provide answer selections and free-text explanations. Table 4 highlights a substantial performance gap, with Gemini-3-Pro and GPT-5.1 achieving accuracies of 45.50% and 39.00%, compared to a human baseline of 74.40%.
Table 4: Accuracy and LLM-judge scores (1-5) on reasoning quality (Rel=Relevance, Cohe=Logical Coherence, Depth=Depth of Analysis, Align=Alignment with Human, Conc=Conciseness, Ovrl=Overall).
Model Acc% Rel Align Cohe Depth Conc Ovrl
Human 74.40 - - - - - -
Gemini-3-Pro 45.50 4.66 3.30 4.67 3.49 4.87 4.10
GPT-5.1 39.00 4.58 3.29 4.65 3.26 4.88 4.00
Gemini-2.5-Pro 37.00 4.57 3.26 4.65 3.41 4.89 4.05
Gemini-2.5-Flash 32.32 4.48 3.17 4.55 3.22 4.87 3.95
GPT-4o-mini 29.00 4.45 3.20 4.56 3.12 4.91 3.90
Qwen2.5-VL-7B 24.50 4.00 2.89 4.21 3.05 4.45 3.63
LLM-judge results show a gap between reasoning presentation and social cognition. Models achieve strong structural scores (Logical Coherence, Relevance >4.0) but substantially lower Alignment and Depth (<3.5). We also find that reasoning similarity to human traces is positively associated with answer correctness (Figure 24, Table 10), proving that robust social interaction understanding requires human-like inference rather than purely fluent justifications. Interestingly, while Gemini-3-Pro maintains a lead in analysis depth over GPT-5.1 (+0.23), this does not improve alignment, indicating that an increased depth of reasoning does not inherently guarantee social accuracy.
A primary failure mode involves the inability to identify core social elements. In 40% to 54% of failures, models prioritize secondary visual cues over the "long-term separation" or "familial bonds" emphasized by humans. Furthermore, 45% of depth-related failures stem from a lack of social dynamic analysis, where models stop at direct causality and ignore power structures or cultural norms. We also observe a unique "excessive caution" bias in GPT-5.1, which frequently refuses to make valid socio-cultural inferences from appearance, resulting in the lowest alignment scores where humans readily use such cues for grounding. These insights suggest that MLLMs require multi-layered social grounding that explicitly partitions inference into observational, affective, and normative layers, enabling the models to bridge the gap between literal perception and human-aligned social intelligence.
@@ 4 Related Works
Video Benchmarks for MLLMs. Rapid progress in video MLLMs has motivated a growing suite of benchmarks spanning perception to high-level reasoning. Representative benchmarks cover multiple tasks (e.g., QA and summarization) over diverse video sources (Zhou et al., 2024; Fu et al., 2024; Li et al., 2024b; Fang et al., 2024; Ning et al., 2023; Wu et al., 2024b). A major line of work emphasizes long-form understanding and temporal reasoning (Rawal et al., 2024; Mangalam et al., 2023; Wang et al., 2024b; Liu et al., 2024). Meanwhile, emerging benchmarks probe new deployment regimes, such as real-time/streaming video reasoning (Xun et al., 2025) and egocentric intent understanding with gaze cues (Peng et al., 2025). While these benchmarks provide broad coverage of video comprehension, they are not designed to systematically evaluate the fine-grained understanding and reasoning required for complex multi-person social interactions.
Evaluating Social Intelligence in AI Systems. Beyond general video understanding, a parallel line of work evaluates social intelligence, moving from component-level signals (e.g., actions and affect) toward richer interpersonal reasoning and grounded social cognition. Recent benchmarks provide narrower but deeper probes, including Theory-of-Mind understanding from short films (Villa-Cueva et al., 2025), grounded social reasoning traces from interaction videos (Mathur et al., 2025; Niu et al., 2026), and nonverbal social reasoning without speech (Li et al., 2025b). Complementary efforts assess emotional and social intelligence in video QA (Zhang et al., 2025) and social intelligence in authentic multi-turn human conversations (Huang et al., 2025), alongside interactive social environments (Zhou et al., 2023; Wang et al., 2024a; Kong et al., 2024). However, existing benchmarks largely evaluate social understanding in a task-specific or modality-specific manner. SIV-Bench complements them by providing a unified evaluation of social interaction understanding in real-world videos, spanning perception, social state reasoning, and dynamics prediction.